OpenStudy (anonymous):

Put the equation in the form ax^2+bx+c=0. (2x+1)^2-3x=5

OpenStudy (anonymous):

expand, combine like terms, subtract 5 form both sides and you will have it

OpenStudy (anonymous):

can write it if you like but it is mainly a matter of multiplying \[(2x+1)(2x+1)\] then subtracting 3x from the result, and then subtracting 5

OpenStudy (anonymous):

morning satellite!

OpenStudy (amistre64):

almost a sensei I see ;)

OpenStudy (anonymous):

morning joe and sensei too!

OpenStudy (anonymous):

no it is too early for the sensei. that comes later in the evening

OpenStudy (amistre64):

12819 exps each level once you get there :)

OpenStudy (anonymous):

whoa as i live and breathe ! who knew? now if i could only be a green owl...

OpenStudy (anonymous):

?? what means 12819 exps?

OpenStudy (anonymous):

(2x+1)^2-3x=5 4x^2+4x+1-3x=5 4x^2+x+1-5=0 4x^2+x-4=0 a=4,b=1,c=-4

OpenStudy (amistre64):

look in your level requirements up top; you need 12819 exp points per level up now

OpenStudy (anonymous):

now sure how many i needed before. a round number maybe? in any case i am waiting for you to see what comes next, because you are the only pioneer. but i think once you get to level 100 you should be able to choose your own honorific